OpenStudy (anonymous):

Which expression shows a complete factorization of x^2(x+1)-9(x+1) ???? I NEED HELP! Thanks! :-)

OpenStudy (anonymous):

just take x+1 common it becomes (x^2-9)(x+1)

OpenStudy (anonymous):

then x^2-9 = x^2-3^3

OpenStudy (anonymous):

that's like a^2-b^2 which becomes (x-3)(x+3)

OpenStudy (anonymous):

so there you get (x+1)(x-3)(x+3)

OpenStudy (anonymous):

P.S. \[a ^{2}-b ^{2} = (a+b) (a-b)\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

you are welcome :)

OpenStudy (raden):

u mean that is x^2-3^2, right :)

OpenStudy (anonymous):

oh yeah typing mistake @RadEn

OpenStudy (raden):

:)